OpenStudy (anonymous):

3f+19=4 What does f equal?

TheSmartOne (thesmartone):

Welcome to OS!!!

TheSmartOne (thesmartone):

All you have to do it subtract 19 and then divide by 3...

TheSmartOne (thesmartone):

3f+19=4 -19 -19 3f=-15 -- --- 3 3 f=-5

TheSmartOne (thesmartone):

@Camille16920

TheSmartOne (thesmartone):

@Camille16920

OpenStudy (anonymous):

f=-5